A student adds two vectors with magnitudes of 180 and 60. Which of the following is NOT a possible magnitude of the resultant?

a. 240

b. 100

c. 180

d. 120

e. 200
Answer #1

Correct Option: b

When two vectors A and B are added, the maximum value of the resultant can be

Mar = A + B

It correspond to the condition when both vectors vectors point in the same direction.

The minimum value is

Min = A - B

It correspond to the condition when both vectors points in the opposite directions.

When two vectors of magnitude 180 and 60 are added, the maximum value can be

Max = 180 + 60 = 240

The minimum value can be

Min = 180 - 60 = 120

All options are possible except b. 100